What is the greatest of 5 consecutive integers if the sum of these integers equals 185?
Answer by alicealc(293)

1st integer = k
2nd integer = k + 1
3rd integer = k + 2
4th integer = k + 3
5th integer = k + 4

k + k+1 + k+2 + k+3 + k+4 = 185
5k + 10 = 185
5k = 185 - 10
5k = 175
k = 175/5
k = 35

so, the greatest integer is k + 4 = 35 + 4 = 39